# Near Field Localization via AI-Aided Subspace Methods

### v1

@@ I Introduction
  • AI-Aided Near-Field MUSIC: We propose a DNN augmentation of 2D near-field MUSIC termed NF-SubspaceNet. We follow the methodology of [49], learning to produce a surrogate covariance that is useful for 2D MUSIC recovery. We overcome the non-differentiability of MUSIC by introducing a loss term inspired by [50], that monitors the spectrum rather than the recovered angles.
@@ III-B2 Architecture
for each $m\in\{1,\ldots,\hat{M}({\boldsymbol{X}};{\boldsymbol{\psi}}_{\rm a})\}$. The $\arg\max$ operation in (18) is computed by grid search over a range holding (1). Sources at the far edge of the range, i.e., at the limit, are labeled as far-field sources. The resulting DCD-MUSIC algorithm is summarized as Algorithm 3.
@@ IV-A2 Evaluation Metrics
We evaluate the RMSPE under varying key parameters, including the SNR, the number of snapshots and the number of sources. Further, we assess robustness in the presence of coherent sources, and array miscalibration, where the exact locations of the sensor elements are unknown to the tested algorithms. In addition to RMSPE, we also assess the algorithms’ ability to estimate the number of sources using the different suggested techniques. The interpretability of our solutions is analyzed through MUSIC spectrum and the beampattern.
(a) Non-coherent sources.
(b) Coherent sources.
(c) Coherent sources & $T=10$.
Figure 4: RMSPE vs. SNR, $M=2$ sources.
@@ IV-A3 Localization Algorithms
We implement NF-SubspaceNet and DCD-MUSIC using the two different configurations for the CNN autoencoder architecture: the default variant uses $3$ encoding and decoding convolutional layers, while the larger variant termed $V2$ employs $4$ encoding and decoding convolutional layers with skip connections.
We compare our results to different model-based algorithms, including Near-Field MUSIC (termed 2D MUSIC), which employs spatial smoothing (SPS) for the coherent case [71]; and standard beamforming  [59]. For non-coherent cases, we also report the Conditional Cramér–Rao Bound (CCRB), using its derivation for non-coherent sources in [72]. As a data-driven benchmark, we use the Transformer MUSIC (TransMUSIC) architecture [40], adapting it to be suitable for the near-field scenario. In settings with unknown number of sources, we used AIC, MDL, or thresholding to estimate $M$.
All methods involving grid searching use the same resolution of $0.5^{\circ}$ in $\big{[}-60^{\circ},60^{\circ}\big{]}$ and $0.5$ meters between the Fresnel and half the Fraunhofer limits. Unless stated otherwise, the training dataset comprises $|\mathcal{D}|=4096$ samples, and the algorithms were evaluated on a test dataset with 1000 different source locations over the specified possible values above. The full details of our architecture and training hyperparameters can be found in our GitHub repository at https://github.com/ShlezingerLab/AI-Subspace-Methods.
@@ IV-B1 Fixed Number of Sources
Our first evaluation compares the considered localization algorithms in the case where the number of sources is fixed at $M=2$. The RMSPE results are presented in Fig. 4 for non-coherent and coherent sources, as well as for a scenario with coherent sources and a short trajectory of only $T=10$ samples. Fig. 4 indicates that when conditions C1 and C3 do not hold, 2D-MUSIC performs well, similar to data-driven approaches. However, when dealing with coherent sources or an insufficient number of snapshots, the model-based algorithm fails to localize accurately. In contrast, both NF-SubspaceNet  and DCD-MUSIC demonstrate robust performance under these challenging conditions. Additionally, due to its high parameterization, TransMUSIC struggles to generalize when trained with only $4096$ samples.
@@ IV-B2 Mixed Number of Sources
We proceed to a more difficult setting in which the number of sources $M$ varies between 2 and 7 sources, drawn uniformly for each sample in the training and test datasets. Here, we compare our NF-SubspaceNet  and DCD-MUSIC with both the default architecture as well as the more parameterized variant $V2$. Additionally, we evaluated different regularization methods to determine which provides better localization and model order estimation for such scenarios. As a baseline, we also included a model-based method for model-order estimation. The results are presented in Table II. A key observation from the results is that increasing the model capacity by adding two more layers consistently improves the model’s ability to localize sources. Furthermore, the V2 models generally demonstrate superior performance in estimating the number of sources. Note that the results of TransMUSIC, which uses an additional DNN for estimating $M$, were not included in the table due to observing a major reduction with localization performance when training the source estimation branch as suggests in [40]. We also observe that DCD-MUSIC combined with AIC typically achieves the most accurate results in estimating the number of sources. Moreover, the regularization introduced to facilitate model-order selection does not necessarily degrade localization accuracy, particularly in more challenging settings with coherent sources, possibly owing to its further encouraging the surrogate covariance to be suitable to downstream subspace-based processing.
Scenario Method Model-Based Acc. NF-SubspaceNet NF-SubspaceNet V2 DCD-MUSIC DCD-MUSIC V2 RMSPE [m] Acc. RMSPE [m] Acc. RMSPE [m] Acc. RMSPE [m] Acc. Non-coherent None - 3.381 - 2.886 - 4.232 - 3.37 - MDL 100 % 3.49 97 % 2.905 85 % 4.616 0 % 3.398 95.1 % AIC 89 % 3.567 89.9 % 3.09 78.3 % 4.793 28 % 3.31 98.1 % Threshold 39 % 3.4 66.9 % 2.9 66 % 4.223 73 % 3.497 67.9 % Coherent None - 8.325 - 7.195 - 9.144 - 7.09 - MDL 0 % 8.251 23.2 % 7.048 66.8 % 9.801 39% 7.07 42.1 % AIC 0 % 8.172 41.7 % 7.425 63.4 % 9.026 58 % 7.163 72.8 % Threshold 0 % 8.022 40.3 % 7.167 45.4 % 8.974 41 % 6.13 48.8 %
@@
The ability of algorithms to simultaneously localize and identify the number of sources is further evident when evaluating these metrics in both non-coherent and fully coherent scenarios for different SNR values. In this scenario, the data size is $|\mathcal{D}|=40,000$. The results are reported in Fig. 5 and Fig. 6. Based on the findings in Table II, we use AIC as the model order estimation method as well as the regularization term in the training stage. The results suggest that for non-coherent case, both data-driven and model-based methods successfully localize up to 7 sources, with the best performance in terms of both RMSPE and localization accuracy achieved using using NF-SubspaceNet. For the coherent case, the data-driven algorithms outperform the model-based ones, and the lowest RMSPE achieved by the TransMUSIC at $\text{SNR}=10\text{dB}$ but with only a minor gap from the two versions of NF-SubspaceNet and DCD-MUSIC V2, which were able to correctly localize the sources as well estimating the number of sources with around 70% .When comparing the performances of the ”small” versions to the V2 versions, we can see that increasing the capacity has better results on both NF-SubspaceNet  and DCD-MUSIC in terms of both localization and model-order estimation accuracy.
@@ IV-B3 Miscalibarion and Limited Snapshots
We next consider challenging settings induced by non-calibrated arrays (C2) and few snapshots (C3). The former is simulated by adding noise distributed uniformly in $[-\eta,\eta]$ to each sensor location. The simulation results for these scenarios with $M=3$ sources at 0 dB SNR are reported in Fig. 7. Observing Fig. 7, we note that our algorithms successfully cope with miscalibrated array, with NF-SubspaceNet consistently achieving the most accurate localization. We can also see that the ability of NF-SubspaceNet to overcome insufficient number of snapshot, only $T=10$. In these challenging settings with limited data, TransMUSIC was unable to properly train, and thus its results are not included in the figure.
@@
Sources Range 2D-MUSIC NF-SubspaceNet DCD-MUSIC $\text{Fresnel}:0.5\cdot\text{Fraunhofer}$ ${\bf 1.234\cdot 10^{-4}}$ $2.158\cdot 10^{-2}$ $1.347\cdot 10^{-2}$ (Fresnel approx.) $\text{Fresnel}:0.5\cdot\text{Fraunhofer}$ $5.395\cdot 10^{-2}$ ${4.154\cdot 10^{-2}}$ ${\bf 1.66\cdot 10^{-2}}$ $\text{Fresnel}:1\cdot\text{Fraunhofer}$ $1.124\cdot 10^{-1}$ $1.739\cdot 10^{-1}$ ${\bf 1.174\cdot 10^{-2}}$ $\text{Fresnel}:2\cdot\text{Fraunhofer}$ $2.485\cdot 10^{-1}$ $2.75\cdot 10^{-1}$ ${\bf 1.045\cdot 10^{-2}}$ $\text{Fresnel}:5\cdot\text{Fraunhofer}$ $3.709\cdot 10^{-1}$ $3.811\cdot 10^{-1}$ ${\bf 9.583\cdot 10^{-3}}$ $\text{Fresnel}:10\cdot\text{Fraunhofer}$ $3.753\cdot 10^{-1}$ $3.708\cdot 10^{-1}$ ${\bf 9.337\cdot 10^{-3}}$
@@ IV-B4 Mixed Near-Field/Far-Field Sources
We compare DCD-MUSIC and NF-SubspaceNet with 2D-MUSIC. All methods employ near-field approximations as in (3). The scenario consists of $M=2$ non-coherent sources with an SNR of $10$ dB. The AI-aided models were trained on data simulated using the Fresnel approximation rather than the full steering vector model. Additionally, the search grid for the range was constrained to values within the Fraunhofer limit.
@@ IV-B5 Large-Scale ULA
A key motivation for developing near-field localization algorithms arises due to the anticipated deployment of large antenna arrays [3]. To demonstrate the practical applicability of our method, we simulated a large-scale ULA with $N=64$ elements and a carrier frequency of 5 GHz. In this scenario we tested the algorithms’ ability to cope with calibration error (modeled as a randomized shifts in each sensor location) with SNR of 10 dB. Here, the Fresnel and Fraunhofer distances, calculated using the array diameter as per (1), result in a Fresnel distance of about 3 meters and Fraunhofer limit of about 120 meters. As this changes the range values, we use a larger dataset of size $|\mathcal{D}|=10^{5}$ for training.
The results for this scenario are presented in Fig. 8. For a calibrated array ($\eta=0$), both the model-based algorithm and NF-SubspaceNet  demonstrate good localization performance. In contrast, DCD-MUSIC, which projects a near-field signal into a far field surrogate covariance, struggles to accurately localize due to the dominant influence of errors originating from the angle estimation branch. Under calibration errors, the model-based algorithm degrades significantly, while NF-SubspaceNet  maintains relatively stable performance. Moreover, DCD-MUSIC, as a data-driven approach, is able to outperform the classical model-based methods under these challenging conditions.
Figure 8: Large-scale ULA, 2 non-coherent sources.

### v2

@@ I Introduction
  • AI-Aided Near-Field MUSIC: We propose a deep learning-aided extension of 2D near-field MUSIC, termed NF-SubspaceNet. Our method follows the general methodology of SubSpaceNet [49], which learns to produce a surrogate covariance matrix tailored for subspace-based localization. However, SubSpaceNet relies on the differentiability of far-field estimators such as Root-MUSIC during training, which are not applicable in the near-field setting. To enable end-to-end training in this more challenging regime, we introduce a differentiable loss function inspired by [50], which supervises the MUSIC spectrum itself rather than the final estimated parameters, thus avoiding the need for differentiable post-processing.
@@ III-A2 Architecture
The last layer of the architecture is a learnable skip connection with a learnable parameter $\alpha$, which is constrained to the range $(0,1)$. This parameter introduces a correction term to the empirical covariance of the input signal, ${\boldsymbol{R}}_{x}[\tau=0]$, by defining the surrogate covariance as follows:
${\boldsymbol{\hat{R}}}({\boldsymbol{X}};{\boldsymbol{\psi}})={\boldsymbol{\hat{R}}}({\boldsymbol{X}};{\boldsymbol{\psi}})+\alpha{\boldsymbol{R}}_{x}[\tau=0].$
@@ III-B2 Architecture
for each $m\in\{1,\ldots,\hat{M}({\boldsymbol{X}};{\boldsymbol{\psi}}_{\rm a})\}$. The $\arg\max$ operation in (18) is computed by grid search over a range holding (1). Sources at the far edge of the range, i.e., at the limit, are labeled as far-field sources. The resulting DCD-MUSIC algorithm is summarized as Algorithm 3. The chosen configuration employs ESPRIT for angle estimation followed by MUSIC for range recovery. While alternative designs are possible, such as using Root-MUSIC for angle estimation combined with a spectrum-based loss for range, instead of approximation of the peak finding operation, we opt the ESPRIT-MUSIC configuration with maskpeak based on empirical trials across our evaluated scenarios.
@@ IV Experimental Study
(a) Non-coherent sources.
(b) Coherent sources.
(c) Coherent sources & $T=10$.
Figure 4: RMSPE vs. SNR, $M=2$ sources.
@@ IV-A2 Evaluation Metrics
We evaluate the RMSPE of the position error, which is therefore reported in meters, under varying key parameters, including the SNR, the number of snapshots and the number of sources. Further, we assess robustness in the presence of coherent sources, and array miscalibration, where the exact locations of the sensor elements are unknown to the tested algorithms. In addition to RMSPE, we also assess the algorithms’ ability to estimate the number of sources using the different suggested techniques. The interpretability of our solutions is analyzed through MUSIC spectrum and the beampattern.
Scenario Method 2D-MUSIC NF-SubspaceNet DCD-MUSIC Acc.[%] RMSPE[m] Acc.[%] RMSPE[m] Acc.[%] RMSPE[m] Non Coherent None - 0.674 - 0.716 - 4.578 MDL 100 99.5 0.712 88.6 4.654 AIC 92.2 94 0.723 92.5 5.349 Threshold 32.7 34.7 0.72 40.2 5.5 Coherent None - 13.758 - 7.665 - 10.463 MDL 0 32 7.832 12.1 10.231 AIC 0 58.5 7.303 48.5 10.231 Threshold 0 34.2 7.688 27.2 10.622
@@ IV-A3 Localization Algorithms
We implement NF-SubspaceNet and DCD-MUSIC using a CNN autoencoder architecture with $3$ encoding and decoding convolutional layers. We compare our results to different model-based algorithms, including Near-Field MUSIC (termed 2D MUSIC), which employs spatial smoothing (SPS) for the coherent case [71]; and standard beamforming  [59]. For non-coherent cases, we also report the Conditional Cramér–Rao Bound (CCRB), using its derivation for non-coherent sources in [72]. As for data-driven benchmarks, we use the Transformer MUSIC (TransMUSIC) architecture [40] and the CNN based model suggested [73] (which we coin DeepCNN), adapting them to be suitable for the near-field scenario. In settings with unknown number of sources, we used AIC, MDL, or thresholding to estimate $M$.
All methods involving grid searching use the same resolution of $0.5^{\circ}$ in $\big{[}-60^{\circ},60^{\circ}\big{]}$ and $0.5$ meters between the Fresnel and half the Fraunhofer limits. Unless stated otherwise, the training dataset comprises $|\mathcal{D}|=4096$ samples, we focus on a limited training dataset to match a real life system which labeld data is not easily obtained, and the algorithms were evaluated on a test dataset with 410 different source locations over the specified possible values above. The full details of our architecture and training hyperparameters can be found in our GitHub repository at https://github.com/ShlezingerLab/AI-Subspace-Methods.
@@ IV-B1 Fixed Number of Sources
Our first evaluation compares the considered localization algorithms in the case where the number of sources is fixed at $M=2$. The RMSPE results are presented in Fig. 4 for non-coherent and coherent sources, as well as for a scenario with coherent sources and a short trajectory of only $T=10$ samples. Fig. 4 indicates that when conditions C1 and C3 do not hold, 2D-MUSIC performs well, similar to data-driven approaches. However, when dealing with coherent sources or an insufficient number of snapshots, the model-based algorithm fails to localize accurately. In contrast, both NF-SubspaceNet  and DCD-MUSIC demonstrate robust performance under these challenging conditions. Additionally, due to their high parameterization, both TransMUSIC and DeepCNN, struggle to generalize when trained with only $4096$ samples. Specifically, DeepCNN shows superior performances only for low SNR values.
@@ IV-B2 Mixed Number of Sources
We proceed to a more difficult setting in which the number of sources $M$ varies between 2 and 8 sources, drawn uniformly for each sample in the training and test datasets. Additionally, we evaluated different regularization methods to determine which provides better localization and model order estimation for such scenarios. As a baseline, we also included a model-based method for model-order estimation. The results are presented in Table II. A key observation here is that using the model order selection test directly on the surrogate covariance yields similar results to the classic use of empirical covariance when the sources are non-coherent. For the coherent case, we can achieve up to $60~{}\%$ accuracy, where the model-based approaches fail, while also achieving improved RMSPE performance. Note that the results of TransMUSIC, which uses an additional DNN for estimating $M$, were not included in the table due to observing a major reduction with localization performance when training the source estimation branch as suggests in [40].
Moreover, the regularization introduced to facilitate model-order selection does not necessarily degrade localization accuracy, particularly in more challenging settings with coherent sources, possibly owing to its further encouraging the surrogate covariance to be suitable to downstream subspace-based processing. Specifically, when considering the coherent case with AIC, NF-SubspaceNet manages to improve its localization comparing to training without any regularization.
@@
The ability of our algorithms to simultaneously localize and identify the number of sources is further evident when evaluating these metrics in both non-coherent and fully coherent scenarios for different SNR values. In this scenario, the data size is $|\mathcal{D}|=40,000$ and the test dataset includes $4,000$ samples. The results are reported in Fig. 5 and Fig. 6. Based on the findings in Table II, we use MDL and AIC, for the non-coherent and coherent cases respectively for NF-SubspaceNet, and AIC for DCD-MUSIC, as the model order estimation method as well as the regularization term in the training stage. The results suggest that for non-coherent case, both data-driven and model-based methods successfully localize up to 8 sources, with the best performance in terms of both RMSPE and localization accuracy achieved using using NF-SubspaceNet. For the coherent case, the data-driven algorithms outperform the model-based ones, and the lowest RMSPE achieved by the NF-SubspaceNet at $\text{SNR}=20\text{dB}$ but with only a minor gap from TransMUSIC, that was able to localized better than DCD-MUSIC. NF-SubspaceNet and DCD-MUSIC were able to partly estimate the number of sources with up to 60% accuracy.
@@ IV-B3 Miscalibarion and Limited Snapshots
We next consider challenging settings induced by non-calibrated arrays (C2) and few snapshots (C3). The former is simulated by adding noise distributed uniformly in $[-\eta,\eta]$ to each sensor location. The simulation results for these scenarios with $M=3$ sources at 0 dB SNR are reported in Fig. 7. Observing Fig. 7, we note that our algorithms successfully cope with miscalibrated array, with both NF-SubspaceNet and DCD-MUSIC consistently achieving the most accurate localization. We can also see that the ability of NF-SubspaceNet to overcome insufficient number of snapshot, only $T=10$.
@@
Sources Range 2D-MUSIC NF-SubspaceNet DCD-MUSIC $\text{Fresnel}:0.5\cdot\text{Fraunhofer}$ ${\bf 4.527\cdot 10^{-5}}$ $5.281\cdot 10^{-5}$ $1.347\cdot 10^{-2}$ (Fresnel approx.) $\text{Fresnel}:0.5\cdot\text{Fraunhofer}$ $5.395\cdot 10^{-2}$ ${4.154\cdot 10^{-2}}$ ${\bf 1.66\cdot 10^{-2}}$ $\text{Fresnel}:1\cdot\text{Fraunhofer}$ $1.124\cdot 10^{-1}$ $1.739\cdot 10^{-1}$ ${\bf 1.174\cdot 10^{-2}}$ $\text{Fresnel}:2\cdot\text{Fraunhofer}$ $2.485\cdot 10^{-1}$ $2.75\cdot 10^{-1}$ ${\bf 1.045\cdot 10^{-2}}$ $\text{Fresnel}:5\cdot\text{Fraunhofer}$ $3.709\cdot 10^{-1}$ $3.811\cdot 10^{-1}$ ${\bf 9.583\cdot 10^{-3}}$ $\text{Fresnel}:10\cdot\text{Fraunhofer}$ $3.753\cdot 10^{-1}$ $3.708\cdot 10^{-1}$ ${\bf 9.337\cdot 10^{-3}}$
@@
(a) Postion RMSPE
(b) Angle RMSPE
(c) Range RMSPE
Figure 8: RMSPE vs. Calibration error, $M=2$ non-coherent sources.
@@ IV-B4 Mixed Near-Field/Far-Field Sources
We compare DCD-MUSIC and NF-SubspaceNet with 2D-MUSIC. All methods employ near-field approximations as in (3). The scenario consists of $M=2$ non-coherent sources with an SNR of $20$ dB. The AI-aided models were trained on data simulated using the Fresnel approximation rather than the full steering vector model. Additionally, the search grid for the range was constrained to values within the Fraunhofer limit.
@@ IV-B5 Large-Scale ULA
A key motivation for developing near-field localization algorithms arises due to the anticipated deployment of large antenna arrays [3]. To demonstrate the applicability of our method, we simulated a large-scale ULA with $N=64$ elements and a carrier frequency of 5 GHz. In this scenario we tested the algorithms’ ability to cope with calibration error (modeled as a randomized shifts in each sensor location) with SNR of 20 dB. Here, the Fresnel and Fraunhofer distances, calculated using the array diameter as per (1), result in a Fresnel distance of about 3 meters and Fraunhofer limit of about 120 meters. As this changes the range values, we use a larger dataset of size $|\mathcal{D}|=10^{5}$ for training. The data-driven benchmarks failed to learn in such large scale settings from such limited data, and are thus not reported here.
The results for this scenario are presented in Fig. 8. Here, in addition to position results, we also report the individual accuracy in recovering the angle and range, to examine the differences in methods performances. In general, we can see that DCD-MUSIC is able to easily adapt to mis-calibrated array, achieving consistent performances for different strength of calibration error. As expected, both NF-SubspaceNet and MUSIC achieve better results when the array is fully calibrated ($\eta=0$). In addition, we can see that NF-SubspaceNet is most accurate in recovering the position, which likely follows from its capability to correctly estimate the range.
@@ IV-C Interpretability
IV-D Time Complexity
We next complement the complexity analysis of Subsection III-C, and add an empirical time analysis in addition to assessing non-asymptotic complexity by counting Multiply-Accumulate Operationss. The time analysis was performed by averaging the inference time of each model over $500$ trials from the setup of Fig. 4, each passed with a batch size of 1. The MACs analysis was done using the THOP library [^1] https://pypi.org/project/thop/. The simulations were run on the same CPU, without threading, over a Linux OS, and ignoring any potential acceleration that could be achieved by using GPUs. The results are shown in Table IV. We can observe that, although NF-SubspaceNet has fewer MACs operations than DCD-MUSIC, its inference time is higher due to the expensive 2D grid search involved in its architecture. Furthermore, TransMUSIC has a slower inference time despite having fewer MACs operations, likely due to the use of the Transformer architecture, which is known for being more computationally intensive in terms of memory and operation complexity.
Model MACs (M) Inference Time (ms) DCD-MUSIC 29.383 5.691 NF-SubspaceNet 14.692 15.160 DeepCNN 212.774 16.527 TransMUSIC 19.843 18.086
TABLE IV: Latency and MACs comparison